OpenStudy (anonymous):

Convert 180 pounds to grams (there are 2.2 lbs per kilogram) Fill out the blanks Using _____ * ____ = _____

OpenStudy (anonymous):

the _____ * ____ = ____ is not exact, but using that method

OpenStudy (anonymous):

\[180lbs \times \frac{ 1kg }{ 2.2lbs }\times \frac{ 1000g }{ 1kg }\]

OpenStudy (anonymous):

Thanks